# Target Specific De Novo Design of Drug Candidate Molecules with Graph Transformer-based Generative Adversarial Networks

### v6

@@ Abstract
Discovering novel drug candidate molecules is one of the most fundamental and critical steps in drug development. Generative deep learning models, which create synthetic data given a probability distribution, offer a high potential for designing de novo molecules. However, for them to be useful in real-life drug development pipelines, these models should be able to design drug-like and target-centric molecules. In this study, we propose an end-to-end generative system, DrugGEN, for the de novo design of drug candidate molecules that interact with intended target proteins. The proposed method represents molecules as graphs and processes them via a generative adversarial network comprising graph transformer layers. The system is trained using a large dataset of drug-like compounds and target-specific bioactive molecules to design effective inhibitory molecules against the AKT1 protein, which is critically important in developing treatments for various types of cancer. We conducted molecular docking and dynamics to assess the target-centric generation performance of the model, as well as attention score visualisation to examine model interpretability. Results indicate that our de novo molecules have a high potential for interacting with the AKT1 protein at the level of its native ligands. Using the open-access DrugGEN codebase, it is possible to easily train models for other druggable proteins, given a dataset of experimentally known bioactive molecules.
@@ 1 Introduction
To address problems associated with conventional de novo drug design, such as long development durations, elevated costs, and a high number of unknown variables regarding the efficacy and safety of the designed compounds, artificial intelligence (AI) driven methods (e.g., deep generative modelling/AI) are starting to emerge in the field [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. A short review of this topic can be found in the Supplementary Material (S1). Most of the methods developed so far focused on designing valid molecules that are typically small in terms of molecular size (via training the models on datasets such as QM9 ($\sim$ 9 heavy atoms on average) [18, 19] and ZINC250K ($\sim$ 23 heavy atoms on average) [8], which is an easier task compared to designing molecules that have the same size distribution as the existing drugs and bioactive molecules ( 29 heavy atoms in ChEMBL molecules, on average).
Deep generative models have also been used to design molecules with desired properties via conditioning the training or the prediction procedure(s). Most of these models have utilised condition vectors as a tool for property injection into the generative process. VAEs [20, 21, 22], GANs [10, 23, 24], sequence-based (language) models [25, 26, 27, 28], geometric models [29] and diffusion models [15, 30] have been utilised for conditional molecule generation tasks. Reinforcement learning (RL) has also been employed for this purpose, with reward-penalty functions guiding models towards desired molecular characteristics in the respective latent space [27, 31]. These approaches result in optimised molecule generation; however, one of the fundamental objectives of drug design is developing small molecules that will physically interact with the desired target, and obtaining de novo molecules with optimised properties alone is insufficient to satisfy this goal. Although a few recent studies present prototype models [32, 33, 34, 35, 36, 37, 38, 39], AI-driven target-centric drug design is a highly novel and understudied area with great potential to contribute to rational drug design.
Figure 1: The workflow of the study. (A) Preparation of datasets including molecules and bioactivities, together with the graph-based encoding of samples in datasets, (B) the graph transformer GAN-based architecture of DrugGEN, model training and evaluation via subjecting de novo molecules to fundamental benchmarks and drug-likeness-related metrics, and (C) subsequent selection of molecules via a series of in silico experiments to identify promising candidates that effectively target the selected protein, (D) the schematic representation of the architecture of the DrugGEN model. The generator module of the GAN, which consists of a graph transformer encoder module, transforms the given input into a new molecular representation. The MLP-based discriminator of the GAN compares the generated de novo molecules to the known inhibitors of the given target protein, scoring them for their assignment to the classes of “real” and “fake” molecules (abbreviations; MLP: multi-layered perceptron, Norm: normalisation, Concat: concatenation, MatMul: matrix multiplication, ElementMul: element-wise multiplication, Mol. adj: molecule adjacency tensor, Mol. Anno: molecule annotation matrix, Upd: updated).
@@ 2.1 The Overview of DrugGEN
DrugGEN aims to learn the distributions of physicochemical properties and topological attributes of drugs and drug candidate molecules from the given data to generate drug-like small molecules that are valid and novel. The discriminator network takes the de novo molecules designed by the generator and compares them with the known (real) bioactive ligands (inhibitors) of the selected target protein so that the differences (in terms of structural properties) between the de novo and real molecules can be learned by the model. This information is then shared with the generator network in an adversarial setting. This approach directs the generation process toward ligands interacting with the selected target. The main technical novelty in the DrugGEN model is the incorporation of graph transformers into the generator and discriminator modules of the generative adversarial network (GAN) (Figure 1D), the aim of which was to overcome problems related to handling the complexity of the target-centric generation process, especially considering the large sizes of drug-like molecules, by effectively learning the long-range dependencies between the atoms of small molecules, and overall, their interactions.
DrugGEN is trained with small molecule data, including random drug-like small molecules and known target-specific inhibitors (from ChEMBL and DrugBank databases), and applied for designing de novo ligands of the “RAC-alpha serine/threonine-protein kinase” (AKT1) protein. Its effectiveness was assessed through a series of downstream data analyses (Figure 1B and C). The performance of DrugGEN was evaluated via known benchmarking metrics. DrugGEN’s molecules were also explored in comparison to real molecules via (i) physicochemical property value distribution plots and (ii) UMAP / t-SNE projections and visualisation in 2-D. Additionally, target-centric properties of the generated molecules were evaluated further via in silico experiments, such as molecular docking and deep learning-based drug-target interaction prediction. In our tests, we also compared DrugGEN with other generative models. Furthermore, we conducted an explainable AI-centric analysis via attention score visualisation to interpret the model. We selected 30 promising candidate de novo molecules via expert curation. Finally, we performed molecular dynamics simulations with (i) the native ligand in the co-crystal structure of AKT1 (PDB id: “4GV1”) [42], and (ii) a chosen de novo generated molecule to compare their binding characteristics within the AKT1 binding site. Each of these analyses is described in detail in the subsequent sections.
2.2 Performance Evaluation and Comparison
@@
We compared DrugGEN’s performance with methods from the molecular generative AI literature. For this, we generated 10,000 de novo molecules (using our trained model in the inference mode) and subjected these molecules to MOSES benchmarking [43]. Here, we report the generative performance of DrugGEN and other models using the widely known metrics of validity, uniqueness, novelty, internal diversity (IntDiv), and quantitative estimate of drug-likeness (QED) where higher values indicate better performance (see “Performance metrics” in Methods for details). It is important to note that these benchmarking metrics only provide clues about the performance of a generative model on a basic level. They do not offer a comprehensive evaluation, which is a difficult task since the main aims of these models are different from each other (e.g., designing a valid molecule, optimising molecules over a physicochemical property, etc.). Target-specific molecule generation, which is the main goal of DrugGEN, cannot be assessed by these metrics; therefore, this property has been evaluated at later steps via physics-based analyses such as molecular docking and dynamics.
@@
In Table 1, DrugGEN’s performance is shown together with other models from the literature (i.e., ORGAN, MolGPT, MGM, RELATION, REINVENT, MARS, BIMODAL, molDQN, QADD, STAGAN, ResGen and EDM). The competing models were selected based on the underlying algorithms and training datasets to cover a wide range of approaches. Since this analysis has multiple metrics and numerous methods, it was not straightforward to make a generalised comparison. To solve this issue, we calculated an overall ranking by considering all the metrics together (where rank 1 is the best). For this, we divided the methods into three groups: the best, medium and low, and scored them with 1, 2 and 3, respectively, independently for each metric. The methods are ranked from lowest to highest average score. DrugGEN became the top model in the overall ranking.
Considering individual metrics, DrugGEN displayed a high performance on most of them and did not suffer from issues such as low validity, novelty or uniqueness, which is the opposite of ORGAN, BIMODAL, REINVENT, MGM and ResGen. Of these models, ORGAN relies on the GAN structure composed of an RNN (as a generator) and a CNN (as a discriminator) to generate conditioned molecules [44]. Another GAN-based model, STAGAN, was introduced to mitigate mode collapse observed in conventional GANs. Nevertheless, the model exhibits diminished validity, particularly when confronted with the challenge of processing considerably larger molecular structures within the ZINC250k dataset ( 23 heavy atoms on average) [11]. On the other hand, DrugGEN utilises graph transformers with attention modules, which might contribute to the performance difference.
@@
Table 1: Molecule generation performance of DrugGEN and other methods: RELATION [38],MGM [12], MolGPT [26], ORGAN [44], QADD [45], molDQN [46], BIMODAL [47], MARS [48], REINVENT [49], ResGen [29], EDM [15], and STAGAN [11] were given in terms of fundamental benchmarking metrics of internal diversity (IntDiv), novelty, validity and uniqueness (Uniq.). We directly obtained the scores from the articles of these models, except for the EDM, for which we re-trained the model with the ChEMBL dataset, and ResGen, for which we used the trained model to generate molecules and calculate metrics. The ORGAN and STAGAN models are trained on a curated ZINC dataset, whereas the RELATION and MolDQN models utilise a combination of ChEMBL and ZINC data for training. The remaining models are trained exclusively on the ChEMBL data. Arrows show the direction in which the performance increases. The best performances are displayed in bold font (all within %1 of the highest score).
@@
Data Type Model Name Architecture details Validity ($\uparrow$) Uniqueness ($\uparrow$) Novelty ($\uparrow$) IntDiv ($\uparrow$) QED ($\uparrow$) Overall ranking ($\downarrow$) Text REINVENT RNN + RL 0.940 0.307 - 0.755 0.525 6 BIMODAL RNN 0.997 0.314 - 0.720 0.541 6 RELATION BiTL 0.854 1.000 1.000 0.773 - 3 MolGPT Transformer 0.994 0.797 1.000 0.857 - 2 ORGAN GAN 0.379 0.687 0.841 - 0.520 8 Graph QADD GNN + RL 1.000 0.341 - 0.613 - 9 MARS GNN + MCMC 0.997 0.333 - 0.641 - 9 MGM MPNN 0.849 0.722 1.000 - 0.582 3 molDQN Q-Learning (RL) 1.000 0.360 - 0.531 - 9 STAGAN GAN 0.482 1.000 1.000 - - 5 ResGen GVP - 1.000 0.289 0.547 0.218 12 DrugGEN (ours) GAN + Graph transformer 0.931 0.991 1.000 0.875 0.520 1 *RNN: Recurrent Neural Networks, RL: Reinforcement Learning, BiTL: Bidirectional Transfer Learning, GAN: Generative Adversarial Networks, GNN: Graph Neural Networks, MCMC: Markov chain Monte Carlo, MPNN: Message-passing Neural Networks, GVP: Geometric Vector Perceptron.
The MolGPT model employs the conventional sequence-based transformer architecture utilised in natural language processing. This model was trained using SMILES representations of molecules. However, the novelty score of the model was reported to be subpar. Furthermore, models such as QADD, molDQN, BIMODAL, REINVENT, and MARS also exhibit low novelty scores caused by generating existing samples in the training set. On the contrary, DrugGEN generates a significantly higher rate of novel molecules (Table 1). We also calculated the average maximum and mean pairwise Tanimoto-based similarity values between DrugGEN’s de novo molecules and all real molecules in the ChEMBL dataset and found 55% and 11% similarity, respectively, which points to the high originality of DrugGEN molecules. The internal diversity (IntDiv) metric indicates the diversity of structures among generated samples. DrugGEN also exhibited notably high performance in IntDiv, demonstrating its ability to learn different molecular structures from the training dataset. Additional comments can be found in Supplementary Material (S3).

### v7

@@ Abstract
Discovering novel drug candidate molecules is one of the most fundamental and critical steps in drug development. Generative deep learning models, which create synthetic data given a probability distribution, offer a high potential for designing de novo molecules. However, to be utilisable in real-life drug development pipelines, these models should be able to design drug-like and target-centric molecules. In this study, we propose an end-to-end generative system, DrugGEN, for the de novo design of drug candidate molecules that interact with intended target proteins. The proposed method represents molecules as graphs and processes them via a generative adversarial network comprising graph transformer layers. The system is trained using a large dataset of drug-like compounds and target-specific bioactive molecules to design effective inhibitory molecules against the AKT1 protein, which is critically important in developing treatments for various types of cancer. We conducted molecular docking and dynamics to assess the target-centric generation performance of the model, as well as attention score visualisation to examine model interpretability. In parallel, selected compounds were chemically synthesised and evaluated in the context of in vitro enzymatic assays, which identified two bioactive molecules that inhibited AKT1 at low micromolar concentrations. These results indicate that DrugGEN’s de novo molecules have a high potential for interacting with the AKT1 protein at the level of its native ligands. Using the open-access DrugGEN codebase, it is possible to easily train models for other druggable proteins, given a dataset of experimentally known bioactive molecules.
@@ 1 Introduction
To address problems associated with conventional de novo drug design, such as long development durations, elevated costs, and a high number of unknown variables regarding the efficacy and safety of the designed compounds, artificial intelligence (AI) driven methods (e.g., deep generative modelling/AI) are starting to emerge in the field [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. A short review of this topic can be found in the Supplementary Material (S1).
Deep generative models have also been used to design molecules with desired properties via conditioning the training or the prediction procedure(s). Most of these models have utilised condition vectors as a tool for property injection into the generative process. VAEs [18, 19, 20], GANs [10, 21, 22], sequence-based (language) models [23, 24, 25, 26, 27], geometric models (e.g., 3D molecule generation) [28, 29, 16] and diffusion models [15, 30] have been utilised for conditional molecule generation tasks. Reinforcement learning (RL) has also been employed for this purpose, with reward-penalty functions guiding models towards desired molecular characteristics in the respective latent space [25, 31, 32]. These approaches result in optimised molecule generation; however, one of the fundamental objectives of drug design is developing small molecules that will physically interact with the desired target, and obtaining de novo molecules with optimised properties alone is insufficient to satisfy this goal. Although a few recent studies present prototype models [33, 34, 35, 36, 37, 38, 39, 40], AI-driven target-centric drug design is a highly novel and understudied area with great potential to contribute to rational drug design.
@@
Figure 1: The workflow of the study. (A) Preparation of datasets including molecules and bioactivities, together with the graph-based encoding of samples in datasets, (B) the graph transformer GAN-based architecture of DrugGEN, model training and evaluation via subjecting de novo molecules to fundamental benchmarks and drug-likeness-related metrics, and (C) subsequent selection of molecules via a series of in silico experiments to identify promising candidates that effectively target the selected protein, (D) the schematic representation of the architecture of the DrugGEN model with powerful graph transformer encoder modules in both generator and discriminator networks. The generator module transforms the given input into a new molecular representation. The discriminator compares the generated de novo molecules to the known inhibitors of the given target protein, scoring them for their assignment to the groups of “real” and “fake” molecules (abbreviations; MLP: multi-layer perceptron, Norm: normalisation, Concat: concatenation, MatMul: matrix multiplication, ElementMul: element-wise multiplication, Upd: updated, N: total # of heavy atoms, T: # of atom types, B: # of bond types, D: hidden dimension size).
@@ 2.1 The Overview of DrugGEN
DrugGEN aims to learn the distributions of the physicochemical properties and topological attributes of drugs and drug candidate compounds from the given data to generate drug-like small molecules that are valid and novel. The discriminator network takes the de novo molecules designed by the generator and compares them with the known (real) bioactive ligands (inhibitors) of the selected target protein so that the structural differences between the de novo and real molecules can be learned by the model. This information is then shared with the generator network in an adversarial setting. This approach directs the generation process toward ligands interacting with the selected target. One of the technical novelties in DrugGEN is the incorporation of graph transformers into the generator and discriminator modules of the generative adversarial network (GAN) (Figure 1D). The reason behind employing graph transformers, as opposed to conventional architectures, e.g., the multi-layer perceptron (MLP), is to overcome problems related to handling the complexity of the target-centric generation process, especially considering the large sizes of drug-like molecules, by effectively learning the long-range dependencies and interactions between atoms/bonds. Graph transformers alone have been used in the literature for molecular property prediction [43, 44, 45] however, to the best of our knowledge, this is the first study to employ molecular graph transformers within a GAN pipeline. Furthermore, we modified the attention mechanism of the graph transformer in both the generator and discriminator modules by enhancing the incorporation of edge information (i.e., atomic bonds), thereby increasing the influence of graph connectivity during molecular learning (please see Methods for details). Finally, in traditional GANs, the input is typically random noise; however, in our model, as in some studies reported in the literature (e.g., CycleGAN [46], and DiscoGAN [47]), real samples are used as input. Due to incorporating an adversarial training framework, we have characterised the architecture of our model as a GAN derivative.
[…]
2.2 Overall performance evaluation and comparison
@@
We compared DrugGEN’s performance with methods from the molecular generative AI literature. For this, we generated 10,000 de novo molecules (using our trained model in the inference mode) and subjected these molecules to MOSES benchmarking [49]. Here, we report the generative performance of models using the widely known metrics of validity, uniqueness, novelty, internal diversity (IntDiv), synthetic accessibility (SA) and quantitative estimate of drug-likeness (QED) (see “Performance metrics” in Methods for details). It is important to note that these benchmarking metrics only provide clues about the performance of a generative model on a fundamental level. They do not offer a comprehensive evaluation, which is a difficult task since the main aims of these models are different from each other (e.g., designing a valid molecule, optimising molecules over a physicochemical property, etc.). Target-specific molecule generation, which is the primary goal of DrugGEN, cannot be assessed by these metrics; therefore, this property has been evaluated at later steps via ML-based bioactivity prediction and physics-based analyses, such as molecular docking and dynamics.
@@
In Table 1, DrugGEN’s performance is shown together with other models from the literature, categorised based on their objectives: non-targeted design models (e.g., REINVENT, BIMODAL, MolGPT, ORGAN, CMGN, EDM, QADD, MARS, MGM, molDQN, and STAGAN) and targeted design models (e.g., RELATION, TRIOMPHE-BOA, ResGen, TargetDiff, and Pocket2Mol). The competing models were selected based on the underlying algorithms and training datasets to cover a wide range of approaches. Here, we report the original results from the published results (unless stated otherwise) due to resource and time constraints and the fact that re-training could risk deviating from the model’s optimised performance. Nevertheless, we retrained the equivariant diffusion model (EDM) [15] using the ChEMBL dataset since its original training data differed. Table 1 includes baseline metrics commonly used for evaluating generative models. Targeted design models were evaluated based on AKT1-specific molecule generation. Since this analysis has multiple metrics and numerous methods, it was not straightforward to make a generalised comparison. To solve this issue, we calculated an overall ranking by considering all the metrics together (where rank 1 is the best). For this, we divided the methods into three groups: the best, medium and low, and scored them with 1, 2 and 3, respectively, independently for each metric. The methods are ranked from lowest to highest average score. The underlying continuous metric values are considered to resolve ties in the score-based ranking. DrugGEN and DrugGEN-NoTarget became the top models in the targeted and non-targeted model rankings, respectively. Additional comments can be found in Supplementary Material (S3).
@@
Table 1: Molecule generation performance of DrugGEN and other methods: RELATION [39],MGM [12], MolGPT [24], ORGAN [50], QADD [51], molDQN [32], BIMODAL [52], MARS [53], REINVENT [54], ResGen [28], CMGN [27],TRIOMPHE-BOA [55], TargetDiff [29], Pocket2Mol [16], and STAGAN [11], measured in terms of fundamental benchmarking metrics of validity, novelty, uniqueness (Uniq.), internal diversity (IntDiv), QED and the overall rank of the models, based on all given metrics. We directly obtained the scores of competing models from their respective articles, except for ResGen, for which we used the trained model to generate molecules and calculate metrics, and EDM, for which we utilised the ChEMBL dataset to train a new model. Performance values missing from the original studies and uncomputable by us are indicated by “-”. The ORGAN and STAGAN models are trained on a curated ZINC dataset, whereas the RELATION and MolDQN models utilise a combination of ChEMBL and ZINC data for training. The remaining models are trained exclusively on the ChEMBL data. Arrows show the direction in which the performance increases. The best performances are displayed in bold font, separately for non-targeted and targeted design model groups (all scores within 1% of the highest value are included, as the differences among these close values are negligible). Metrics that contain a distribution of values are reported as mean ± SD, while metrics yielding a single value are reported as such.
@@
Model objective Model Name Input data modality Architecture details * Dataset Validity ($\uparrow$) Novelty ($\uparrow$) Uniq. ($\uparrow$) IntDiv ($\uparrow$) QED ($\uparrow$) Overall ranks ($\uparrow$) Non- targeted design REINVENT Text RNN + RL ChEMBL 0.940 0.307 - 0.755 0.525 10 BIMODAL Text RNN ChEMBL 0.997 0.314 - 0.720 0.541 9 MolGPT Text Transformer MOSES 0.996 0.782 1.000 0.857 - 2 ORGAN Text GAN ZINC 0.379 0.687 0.841 - 0.520 12 CMGN Text Transformer ZINC + ChEMBL 0.989 0.585 0.409 0.682 - 11 EDM 3D coord.* Diffusion models ChEMBL 0.106 1.000 0.964 0.905 0.223 5 QADD Graph GNN + RL ChEMBL 1.000 0.341 - 0.614 ± 0.029 0.785 ± 0.008 6 MARS Graph GNN + MCMC ChEMBL 0.997 0.333 - 0.641 0.746 8 MGM Graph MPNN ChEMBL 0.849 0.722 1.000 - 0.582 4 molDQN Graph Q-Learning (RL) ZINC + ChEMBL 1.000 0.360 - 0.531 0.761 7 STAGAN Graph GAN ZINC-250k 0.482 1.000 1.000 - - 3 DrugGEN- NoTarget (ours) Graph Graph transformer GAN ChEMBL 0.913 0.990 1.000 0.874 ± 0.020 0.502 ± 0.214 1 Targeted design RELATION Text BiTL ChEMBL + PDBBind 0.854 1.000 0.999 0.857 ± 0.016 0.558 ± 0.187 2 TRIOMPHE- BOA Text VAE + Attention ExCAPE - - 0.445 0.586 ± 0.156 0.819 ± 0.116 5 ResGen 3D coord. GVP CrossDocked2020 - - 0.289 0.547 ± 0.092 0.218 ± 0.096 6 TargetDiff 3D coord. SE(3) Diffusion CrossDocked2020 0.839 - 1.000 0.879 ± 0.012 0.462 ± 0.190 3 Pocket2Mol 3D coord. E3GNN + MCMC CrossDocked2020 0.962 - 0.365 0.844 ± 0.022 0.400 ± 0.188 4 DrugGEN (ours) Graph Graph transformer GAN ChEMBL 0.931 0.991 1.000 0.866 ± 0.018 0.507 ± 0.213 1 *RNN: Recurrent Neural Networks, RL: Reinforcement Learning, BiTL: Bidirectional Transfer Learning, GAN: Generative Adversarial Networks, GNN: Graph Neural Networks, MCMC: Markov chain Monte Carlo, MPNN: Message-passing Neural Networks, GVP: Geometric Vector Perceptron. coord.: coordinates
2.3 DrugGEN generates target-specific drug-like molecules
The DrugGEN model was trained and tested on two distinct targets, AKT1 and CDK2, to evaluate its performance across different proteins. The evaluation was mainly conducted for the AKT1 protein, with CDK2 included as an additional target to assess the reproducibility of our system. We selected CDK2 because it is frequently deregulated in various cancers, making it a critical target for de novo drug development [56]. Additionally, there are CDK2 targeting models in the literature that can be compared with DrugGEN.
In this analysis, we evaluated targeted de novo design models, also presented in Table 1, namely RELATION, TRIOMPHE-BOA, ResGen, TargetDiff, Pocket2Mol and DrugGEN, using docking scores against AKT1 and CDK2, together with drug-likeness-related metrics, including quantitative estimate of drug-likeness (QED), synthetic accessibility (SA) [49] Frechet ChemNet Distance (FCD), Lipinski, Veber, and PAINS. We also added fragment and scaffold similarities to compare de novo molecules with real inhibitors of the selected targets (see “Performance metrics” in Methods for details). Finally, we calculated these metrics for the training dataset (real) molecules to serve as a reference. Table 2 presents the results of this comparative analysis. Also, Figure 2A shows docking results with bar graphs (percentages shown on top of each bar represent docking performance of generative models compared to the real AKT1/CDK2 inhibitors, i.e., as the percentage of the docking performance of real inhibitors).
In AKT1-targeted de novo molecule generation, DrugGEN scored the highest docking performance with -8.386 kcal/mol and 99.98% (Table 2 and Figure 2A), exhibiting a noteworthy contrast to the other models. Furthermore, the difference between the DrugGEN and the next best model, RELATION (i.e., -8.386 and -8.100 kcal/mol, respectively), was found to be statistically significant according to the Mann-Whitney U rank test (p-value < 0.0001). Following DrugGEN and RELATION, TargetDiff and Pocket2Mol exhibit satisfactory docking scores (91.59% and -7.981 kcal/mol for TargetDiff and 91.10% and -7.957 kcal/mol for Pocket2Mol). DrugGEN also achieved a well-balanced profile across QED (0.507), SA (3.510), and structural similarity metrics.
When DrugGEN is trained for targeting the CDK2 protein, it achieved a top 10% median docking score of -8.747 kcal/mol (86.99%), indicating superior binding affinity compared to TargetDiff while remaining competitive with Pocket2Mol (Table 2). Our model generated drug-like molecules with an average QED score of 0.547, reflecting its capability to design compounds that adhere to key drug-likeness criteria, further supported by its high performance on the Lipinski, Veber, and PAINS filters (Table 2). On the other hand, Pocket2Mol showed suboptimal drug-likeness characteristics, with lower QED and higher SA scores (Table 2), indicating challenging synthetic complexities. More details on docking and drug-likeness analysis can be found in Supplementary Material (S4).
Furthermore, we measured the docking binding free energy of the capivasertib-AKT1 complex (i.e., the bound ligand in our template PDB structure: 4GV1), as well as other experimentally proven inhibitors that reached clinical trials, namely ipatasertib and uprosertib, at -9.517, -9.681 and -8.569 kcal/mol, respectively. DrugGEN was able to generate 40 molecules that surpass the binding score of the native ligand capivasertib, which are shown in Figure S2.
Table 2: Drug-likeness related and target-centric performance of DrugGEN and other methods: RELATION[39], ResGen [28], TRIOMPHE-BOA [55], TargetDiff [29], and Pocket2Mol [16], measured in terms of of QED, synthetic accessibility (SA), FCD, fragment similarity, scaffold similarity, and adherence to Lipinski, Veber, and PAINS filters, together with docking scores (median kcal/mol values of the top 10% molecules in terms of docking scores), for the AKT1 and CDK2 targeting tasks, separately. Values for the training dataset (real molecules) are also provided. Arrows show the direction in which the performance increases. The best performances are displayed in bold font, separately for AKT1 and CDK2 targeting models (all within %1 of the highest score are included due to very close numbers without a significant difference). Performance values that could not be calculated are indicated by “-”. Metrics that contain a distribution of values are reported as mean ± SD, while metrics yielding a single value are reported as such.
Model objective / data property Model/dataset name Docking score ($\downarrow$) QED ($\uparrow$) SA ($\downarrow$) FCD ($\downarrow$) Fragment Sim. ($\uparrow$) Scaffold Sim. ($\uparrow$) Lipinski ($\uparrow$) Veber ($\uparrow$) PAINS ($\uparrow$) Training Dataset (real molecules) Random ChEMBL mol.s* - 0.537 ± 0.223 3.008 ± 0.965 - - - 4.645 ± 0.801 1.827 ± 0.474 0.935 Real AKT1 inhibitors -8.387 ± 0.048 0.552 ± 0.189 3.225 ± 0.833 - - - 4.736 ± 0.507 1.951 ± 0.236 0.947 Real CDK2 inhibitors -9.352 ± 0.041 0.580 ± 0.148 2.911 ± 0.591 - - - 4.927 ± 0.289 1.954 ± 0.213 0.910 AKT1 targeting RELATION -8.100 ± 0.038 0.558 ± 0.187 2.711 ± 0.546 20.186 0.770 0.267 4.854 ± 0.394 1.961 ± 0.198 0.948 TRIOMPHE-BOA -6.920 ± 0.061 0.819 ± 0.116 2.534 ± 0.497 69.638 0.201 0.000 4.997 ± 0.056 1.976 ± 0.153 0.525 ResGen -6.468 ± 0.101 0.218 ± 0.096 5.680 ± 1.208 42.100 0.185 0.000 3.012 ± 0.941 1.171 ± 0.377 0.491 Pocket2Mol -7.957 ± 0.033 0.400 ± 0.188 4.414 ± 1.134 26.413 0.638 0.000 4.225 ± 0.688 1.74 ± 0.441 0.915 TargetDiff -7.981 ± 0.025 0.462 ± 0.19 4.952 ± 0.857 30.663 0.598 0.003 4.673 ± 0.681 1.724 ± 0.498 0.964 MLP* baseline -6.367 ± 0.017 0.545 ± 0.187 3.117 ± 0.923 36.227 0.519 0.003 4.815 ± 0.496 1.895 ± 0.319 0.966 DrugGEN (ours) -8.386 ± 0.018 0.507 ± 0.213 3.510 ± 0.809 20.416 0.721 0.246 4.599 ± 0.815 1.834 ± 0.437 0.932 CDK2 targeting Pocket2Mol -8.803 ± 0.030 0.387 ± 0.187 3.996 ± 0.902 30.119 0.484 0.000 4.421 ± 0.63 1.614 ± 0.487 0.907 TargetDiff -8.353 ± 0.024 0.556 ± 0.185 5.017 ± 1.008 26.483 0.766 0.000 4.833 ± 0.508 1.888 ± 0.334 0.967 DrugGEN (ours) -8.747 ± 0.022 0.547± 0.180 3.409 ± 0.925 10.611 0.931 0.311 4.860 ± 0.442 1.906 ± 0.314 0.940
*MLP: Multilayer perceptron, mol.s: molecules
Figure 2: Exploration of de novo molecules via downstream analysis. (A) Bar plots displaying the median of binding free energies measured in the docking analysis of de novo molecules generated by DrugGEN, RELATION, TargetDiff, Pocket2Mol, TRIOMPHE-BOA, and ResGen models (molecules are docked into the binding pocket of the AKT1/CDK2 protein structure). The whiskers on the bars represent the standard error of the median (median values are utilised due to the non-normal distribution of data). The percentages above each bar represent docking scores normalised with respect to the scores of real AKT1/CDK2 inhibitors, with the top real AKT1/CDK2 inhibitors set at 100%. (B) 2-D visualisation of the molecules generated by DrugGEN and DrugGEN-NoTarget models, real AKT1 inhibitors and randomly selected ChEMBL molecules via UMAP and t-SNE projections. (C) Deep learning-based bioactivity prediction analysis. Top: the workflow of DEEPScreen, which uses 2-D pixel-based (image) structural representations of molecules as input and processes them via deep convolutional neural networks [57]. Bottom-left: the DEEPScreen AKT1 model is trained with binarised experimental bioactivity data points of the AKT1 protein (obtained from ChEMBL binding assays). Bottom-right: molecules (i.e., DrugGEN and DrugGEN-NoTarget molecules and randomly selected ChEMBL molecules) are submitted to DrugGEN AKT1 DTI prediction model for inference.
Additionally, with the aim of assessing the generative performance of DrugGEN from the perspective of input data, we measured how the type of input molecules (i.e., random ChEMBL molecules vs. real AKT1 inhibitors) given to the model during inference changed the output. The results are provided in Supplementary Material (S5) with 3 example molecules in Figure S3.
Finally, we conducted physicochemical property comparisons between DrugGEN, other target-based generation models, and real molecule datasets (Supplementary Material S7 and Figure S5); and further assessed the drug-related properties of DrugGEN molecules by subjecting them to Lipinski, Veber and PAINS filters (Supplementary Material S8).
Overall, the results demonstrate that the generative and drug-likeness metrics of DrugGEN are consistent across different targets, indicating the versatility of the generation pipeline for diverse protein targets. Therefore, it is possible to state that DrugGEN was successful in its target-specific and drug-like molecular design routine.